# Torus Knots in Adjoint Representation

### v1

@@ Introduction.
Evaluating the HOMFLY-PT invariant of various knots and links is an important goal because of a series of applications. First of all, in accordance with [1], this invariant colored with a Young diagram $R$ is nothing but the Wilson loop average in representation $R$ in the $SU(N)$ Chern-Simons theory. Hence, studying it allows one to study pure topological effects in gauge theories. In particular, the study of adjoint representation gives information about the pure gluodynamics.
The adjoint invariants are also applied to study of the Vogel’s universality of Chern-Simons theory [2] (see also [3] for a latest review), which gives a unified description of various quantities in the adjoint sector of representation theory for all simple Lie algebras in terms of three parameters $\alpha,\beta,\gamma$, which are homogeneous coordinates of the Vogel’s projective plane. These quantities vary from the Chern-Simons partition function [4] to the Racah matrix [5] involving the adjoint representation and its descendants [6]. In particular, in [7, 8] we have constructed universal invariants for torus knots $T[2,2n+1]$, $T[3,3n\pm 1]$, and $T[4,2n+1]$. However, considering more general knots has not been possible so far.
There is also a pure knot theory application: the knowledge of the adjoint invariant allows one to evaluate the fundamental HOMFLY-PT of the reversed 2-cable knot $K$, which has a peculiar “panhandle” shape, in accordance with [9, Eq.(2)]
@@
The simplest class of knots, which we are going to study here is the torus knots $T[m,n]$. They are given by a pair of coprime numbers $m$ and $n$ (the common non-trivial divisor $l$ of $m$ and $n$ would lead to an $l$-component link). More concretely, we calculate the HOMFLY-PT invariant for these knots in the adjoint representation and in the topological framing so that formula (1) looks as
@@
Throughout the paper, we use the notion
$\{x\}:=x-{1\over x},\ \ \ \ \ \ \ \ \ [n]_{q}:={\{q^{n}\}\over\{q\}}.$
@@
is the eigenvalue of the second Casimir operator in the representation $Q$ given by the highest weight vector $\Lambda$,
@@
One of the technically direct ways to evaluate the Adams coefficients is to use the explicit formula
@@
Below we obtain the answer for the adjoint HOMFLY-PT invariant (i.e. for the $A_{N-1}$ series) for an arbitrary torus knot.
$A_{N-1}$ series.
In the $A_{N-1}$ series case, all the ingredients of the Rosso-Jones formula look as follows.
The finite-dimensional irreducible representation is labeled by the Young diagram $R$, and the corresponding character is given by the Schur function $S_{R}$ [13], which is realized either as a graded polynomial of ${\rm tr}\,g^{k}$, where $g$ is the group element of $SL(N)$, or as a graded symmetric polynomial of the eigenvalues of $g$.
@@
We will also need the notion of composite representation [14], which is the most general finite-dimensional irreducible highest weight representation of $SL(N)$, which is associated with the Young diagram of the form
@@
The adjoint representation is the simplest example of the composite representation, it is given by the Young diagram $21^{N-2}]$, which depends on $N$. We introduce the variable $A:=q^{N}$. This variable $A$ should not be confused with the second variable $A$ of the HOMFLY-PT polynomials: the HOMFLY-PT polynomial is evaluated for one and the same Young diagram for various $A_{N-1}$, while the adjoint (or, more generally, composite) representation depends on $N$ itself. Formally, we define $H_{\mathrm{Adj}}(q,A)$ via the specialization
@@
The adjoint quantum dimension and Casimir exponential are
@@
Then, the Adams operation gives [11][12]
@@
$\kappa_{(R,P)}$ $=$ $2Nm+\kappa_{R}-\kappa_{P^{T}}\,$,
@@
Inserting this answer (22) into formula (2), we immediately obtain the fundamental normalized HOMFLY-PT polynomial of the reverse 2-cable $T[m,n]$ knot [9].
@@
With these formulas, one can check for generic torus knot $T[m,n]$ that
@@
  • The special polynomial $\sigma^{T[m,n]}_{\mathrm{Adj}}(A):=H^{T[m,n]}_{\mathrm{Adj}}\Big|_{q=1}$, which, for the series of torus knots $T[2,n=2k+1]$, $T[3,n=3k\pm 1]$, $T[4,n=2k+1]$, $\ldots$, is equal to (the signs in front of $\sigma_{[1]}^{T[m,n]}$ is, of course, not seen in the uniform special polynomial)
@@
  • The $A=q^{2}$ corresponds to the spin 1 representation of $SL(2)$, and, hence, the adjoint HOMFLY-PT polynomial becomes the colored Jones polynomial: $J^{T[m,n]}_{2}(q):=H^{T[m,n]}_{\mathrm{Adj}}\Big|_{A=q^{2}}$. Indeed,
@@
We presented an explicit formula(22) for the adjoint HOMFLY-PT polynomial for an arbitrary torus knot $T[m,n]$. A natural extension of the present analysis is to the case of torus links, which is considerably more involved. The complexity arises for two main reasons. First of all, the number of representations $Q$ that contribute to the r.h.s. of the Rosso-Jones formula in the case of knots is typically less than half that in the case of links. This is because the $m$-plethysm expansion of the adjoint representation gives rise to composite representations $(R,P)$ with $|R|=|P|=m$ only (and not all of possible representations even with this restriction emerge), while, in the case of $l$-component links, the sum in the Rosso-Jones formula runs over representations with $|R|=|P|$ equal to $mk/l$ with $k=1,\ldots,l$ (see [9]). An additional complication arises from the non-trivial multiplicities of representations appearing in the sum. Because of these reasons, calculating the adjoint HOMFLY-PT invariants for arbitrary links remains a future problem to solve.
@@ Acknowledgements.
We are grateful to Alexander Stoimenov and Hisham Sati, whose intelligent questions were one of the reasons that prompted us to conduct this research. The work of A.M. was partially funded within the state assignment of the NRC Kurchatov Institute, was partly supported by the grant of the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” and by Armenian SCS grants 24WS-1C031. The work of V.K.S. is supported by Tamkeen UAE under the NYU Abu Dhabi Research Institute grant CG008

### v2

@@ Introduction.
Evaluating the HOMFLY-PT invariant of various knots and links is an important goal due to its numerous applications. Specifically, in accordance with [1], this invariant colored with a Young diagram $R$ is nothing but the Wilson loop average in representation $R$ in the $SU(N)$ Chern-Simons theory. Hence, its analysis reveals pure topological effects in gauge theories. In particular, the study of adjoint representation gives information about the pure gluodynamics.
The adjoint invariants are also applied to study of the Vogel’s universality of Chern-Simons theory [2] (see also [3] for a latest review), which gives a unified description of various quantities in the adjoint sector of representation theory for all simple Lie algebras in terms of three parameters $\alpha,\beta,\gamma$, which are homogeneous coordinates of the Vogel’s projective plane. These quantities vary from the Chern-Simons partition function [4] to the Racah matrix [5] involving the adjoint representation and its descendants [6]. Notably, in [7, 8], one of us constructed universal invariants for the torus knots $T[2,2n+1]$, $T[3,3n\pm 1]$, and $T[4,2n+1]$. Extending this construction to more general knots, however, has not yet been possible.
Moreover, there is also a pure knot theory application: the knowledge of the adjoint invariant allows one to evaluate the fundamental HOMFLY-PT of the reversed $2$-cable knot $K$, which has a peculiar “panhandle” shape, in accordance with [9, Eq.(2)]
@@
The simplest class of knots that we study here is the torus knots $T[m,n]$. They are given by a pair of coprime numbers $m$ and $n$ (otherwise yielding a multi-component link). More concretely, we calculate the HOMFLY-PT invariant for these knots in the adjoint representation and in the topological framing so that formula (1) looks as
@@
In this paper, we adopt the following notation:
$\{x\}:=x-\frac{1}{x},\qquad[n]_{q}:=\frac{\{q^{n}\}}{\{q\}}.$
@@
is the eigenvalue of the second Casimir operator in the representation $Q$, expressed in terms of its highest weight $\Lambda_{Q}$, and
@@
One of the most straightforward ways to evaluate the Adams coefficients is to use the explicit formula
@@
In the following, we specialize the general formula(3) to the $\textbf{A}_{N-1}$} series and derive explicit expressions for the adjoint HOMFLY-PT invariant of arbitrary torus knots.
$\textbf{A}_{N-1}$} series.
For the $\textbf{A}_{N-1}$} series, all ingredients entering the Rosso–Jones formula (3) admit a particularly explicit form. The finite-dimensional irreducible representation is labeled by the Young diagram $R$, and the corresponding character is given by the Schur function $S_{R}$ [13], which is realized either as a graded polynomial of ${\rm tr}\,g^{k}$, where $g$ is the group element of $SL(N)$, or as a graded symmetric polynomial of the eigenvalues of $g$.
@@
We will also make use of the notion of a composite representation [14], which describes the most general finite-dimensional irreducible highest-weight representation of $SL(N)$. Such a representation is associated with a Young diagram of the form
@@
The adjoint representation is the simplest example of the composite representation, it is given by the Young diagram $21^{N-2}]$, which depends on $N$. We introduce the variable $A:=q^{N}$ as a convenient notation for our discussion. This variable $A$ should not be confused with the second variable $A$ of the HOMFLY-PT polynomials. The HOMFLY-PT polynomial is evaluated for one and the same Young diagram for various $\textbf{A}_{N-1}$}, while the adjoint (or, more generally, composite) representation depends on $N$ itself. Formally, we define $H_{\mathrm{Adj}}(q,A)$ via the specialization
@@
For the adjoint representation, the quantum dimension and the Casimir exponential take particularly simple forms:
@@
Then, the Adams operation gives (see [11],[12])
@@
$\kappa_{(R,P)}$ $=$ $2Nm+\kappa_{R}-\kappa_{P^{T}}\,.$
@@
Substituting the result (22) into (2), we immediately obtain the fundamental normalized HOMFLY-PT polynomial of the reverse 22-cable $T(m,n)$ knot [9]. The following section is devoted to a detailed study of the properties of the obtained formulas(22).
@@
Using the formulas (22), one can verify for a generic torus knot $T(m,n)$ that the following properties hold:
@@
  • The special polynomial
    $\sigma^{T[m,n]}_{\mathrm{Adj}}(A):=H^{T[m,n]}_{\mathrm{Adj}}\Big|_{q=1},$
    which, for the series of torus knots $T[2,n=2k+1]$, $T[3,n=3k\pm 1]$, $T[4,n=2k+1]$, $\ldots$, is equal to (the signs in front of $\sigma_{[1]}^{T[m,n]}$ are, of course, not seen in the uniform special polynomial)
@@
  • The $A=q^{2}$ corresponds to the spin 1 representation of $SL(2)$, and, hence the adjoint HOMFLY-PT polynomial becomes the colored Jones polynomial: $J^{T[m,n]}_{2}(q):=H^{T[m,n]}_{\mathrm{Adj}}\Big|_{A=q^{2}}$. In fact:
@@
We presented an explicit formula(22) for the adjoint HOMFLY-PT polynomial for an arbitrary torus knot $T[m,n]$. A natural extension of the present analysis is to the case of torus links, which is considerably more involved. The complexity arises for two main reasons. First of all, the number of representations $Q$ that contribute to the r.h.s. of the Rosso-Jones formula in the case of knots is typically less than half that in the case of links. This is because the $m$-plethysm expansion of the adjoint representation gives rise to composite representations $(R,P)$ with $|R|=|P|=m$ only (and not all of possible representations even with this restriction emerge). In the case of $l$-component links, the sum in the Rosso-Jones formula runs over representations with $|R|=|P|$ equal to $mk/l$ with $k=1,\ldots,l$ (see [9]). An additional complication arises from the non-trivial multiplicities of representations appearing in the sum. Due to these chain of reasoning, calculating the adjoint HOMFLY-PT invariants for arbitrary links remains a challenging problem. It therefore represents an important topic for future investigation.
@@ Acknowledgements.
We are grateful to Alexander Stoimenov and Hisham Sati, whose intelligent questions were one of the reasons that prompted us to conduct this research. The work of A.M. was partially funded within the state assignment of the NRC Kurchatov Institute, was partly supported by the grant of the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” and by Armenian SCS grants 24WS-1C031. The work of V.K.S. is supported by Tamkeen UAE under the NYU Abu Dhabi Research Institute grant CG008.